Solve the system of equations using the substitution method. IF the answer is a unique solutin, present it as an ordered paird. If not, specifiy whether the answer is no solution or infinitely many solutions and state how you arrived at that conclusion. 3x+y-2, 2x-y=-3
Answer by checkley75(3666)

3x+y(=)2 (I'll assume the - sign should be an = sign) or y=2-3x
2x-y=-3 or y=2x+3
Thus:
2x+3=2-3x
2x+3x=2-3
5x=-1
x=-1/5 ans.
2*-1/5-y=-3
-2/5-y=-3
-y=-3+2/5
-y=-13/5
y=13/5 ans.